8(x-3)-(6-2x)=2(x+2)-5(5-x)

Answer by Cromlix(4381)

8(x-3)-(6-2x)=2(x+2)-5(5-x)
8x - 24 - 6 + 12x = 2x + 4 - 25 + 5x
Collect like terms.
8x + 12x - 2x - 5x = 24 + 6 + 4 - 25
13x = 9
x = 9/13
Hope this helps.
:-)